OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 7 tan3x - 21 tan x = 0

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

same like last one let tan x =a then take out common factor

OpenStudy (anonymous):

so quadratic equation again?

Nnesha (nnesha):

ye... but first you should find out common factor

Nnesha (nnesha):

\[\huge\rm 7a^3 -21 a = 0\] what is common factor ?

OpenStudy (anonymous):

7 is a common factor

Nnesha (nnesha):

yes right and what else ?

OpenStudy (anonymous):

7a(a^2-3)=0

Nnesha (nnesha):

yep right set both 7a and a^2 -3 equal to zero solve for a

OpenStudy (anonymous):

a=0 and a=sqrt.3

Nnesha (nnesha):

yep that means tan x = 0 and tanx = sqrt{3} unit circle!

Nnesha (nnesha):

tan = sin /cos

OpenStudy (anonymous):

so wait what would I look for

Nnesha (nnesha):

tanx = 0 and tan x =sqrt 3 this one z tricky you have to divide sin over cos to find out tan